# Unpacking Tokenization: Evaluating Text Compression and its Correlation with Model Performance

### v1

@@ Abstract
[…]
@@ 2 Measuring Tokenization Quality
From the very early days of NLP, models have always assumed text discretized to tokens as input Winograd (1971). For the most part, these tokens were whitespace separated words, but in the recent decade non-trivial tokenization algorithms have surfaced Mikolov et al. (2012); Sennrich et al. (2016), primarily to deal with unseen tokens without smoothing techniques or other convoluted methods Chen and Goodman (1999). The underlying reasoning behind all modern tokenization methods is that some subwords, e.g., morphemes, may carry independent information that is of value to the model even if the word as a whole is rare or unseen. Therefore, better tokenization is assumed to improve models’ performance over rare words, while also carrying computational benefits, like smaller models and the elimination of unknown tokens.
@@
Other works circle back and assess tokenization with respect to the desiderata it is supposed to serve as a stand-alone algorithm, independently from the model trained on top of it. This is done usually in addition to evaluation over downstream performance. However, most works disagree on the desiderata themselves. Many emphasize alignment to linguistically meaningful units Hofmann et al. (2021, 2022); Klein and Tsarfaty (2020) or to human cognitive preferences in general Beinborn and Pinter (2023). [^1] This line of works may view tokenization as a continuation of unsupervised morphemic segmentation Creutz and Lagus (2002); Virpioja et al. (2013). Others include analyses of token length and frequency Bostrom and Durrett (2020); Yehezkel and Pinter (2023), mostly in addition to the above, assuming that ideal tokenizers use longer and more frequent tokens.
The two types of tokenization evaluations, extrinsic over downstream success and intrinsic over a plethora of metrics, are usually not compared directly. They are only used to demonstrate the superiority of a specific tokenizer, and the relations between the evaluation approaches is glossed over. In this work we explicitly focus on compression as a potential intrinsic indicator of tokenization quality and check to what extent it is correlated with extrinsic downstream success. We conclude that compression is a desideratum for tokenization not only due to its theoretical virtue, expanded on in the next section, but first and foremost because it correlates with downstream performance.
@@ 3 The Role of Compression in Tokenization
But in a sense, compression-driven tokenization may be viewed as language modeling in and of itself. Consider that language modeling is aimed at assessing and possibly minimizing the likelihood of produced texts expressed as a product of token probabilities,
@@
where $x_{k}$ is the $k$th token in a sentence $\mathbf{x}$. Compression limits the upper bound on this product of fractions by minimizing the number of operands, i.e., minimizing the length of the sequence.
@@
$P(\mathbf{x})\approx\prod_{k}P(x_{k}|\mathbf{x}_{k-n:k-1}),$
and the probability is further approximated by the abundance of the relevant $n$-gram in a training corpus,
$P(x_{k}|\mathbf{x}_{k-n:k-1})\propto C(\mathbf{x}_{k-n:k}),$
@@
We set to empirically examine our hypothesis by assessing the correlation between the tokenizer’s compression ability and the performance of language models of various sizes over a set of downstream tasks. To control the compression ability, we recall that BPE’s maximal compression is guaranteed only over its supporting corpus. Normally, for large enough corpora, a minimal discrepancy is assumed between the character distribution in the corpus and the “true” distribution in the language. In this work however, we explicitly emphasize and expand this discrepancy by limiting the size of the support to a great extent. We will show that this intervention severely hinders the compression capacity of the tokenizer and that it also leads to deteriorating downstream performance.
@@ 4.1 English Experiments
[…]
@@ 5.1 Intrinsic Evaluation
To illustrate the effect of limiting the tokenization support on the compression ability, we measured the accumulative length in tokens of the development sets of all English tasks used in downstream.
@@
Table 3: Spearman’s $\rho$ coefficient for rank correlation between the tokenizers’ support and the downstream performance for all model sizes and tasks. Asterisks are used to denote statistically significant correlation.                                                                                                                                   * $p<0.05$, ** $p<0.01$
@@
Table 3: Spearman’s $\rho$ coefficient for rank correlation between the tokenizers’ support and the downstream performance for all model sizes and tasks. Asterisks are used to denote statistically significant correlation.                                                                                                                                   * $p<0.05$, ** $p<0.01$
Table 4: Pearson’s correlation coefficient between the tokenizers’ compression and the downstream performance for all model sizes and tasks. Asterisks are used to denote statistically significant correlation.                                                                                                                                   * $p<0.05$, ** $p<0.01$
@@ 7 Conclusions
In this paper we demonstrated the importance of compression to tokenization, both as an intrinsic evaluation of tokenization quality and as an indicator for better performance on extrinsic downstream tasks. We argued in favor of compression-driven tokenization from a theoretical perspective, since it may make the tokenizer act as a simple standalone language model, and we showed its correlation with downstream model success.

### v2

@@ Abstract
Omer Goldmanβγ    Avi Caciularuγ    Matan Eyalγ Kris Caoδ   Idan Szpektorγ   Reut Tsarfatyγ βBar-Ilan University   γGoogle Research   δGoogle DeepMind {ogoldman,avica,matane,kriscao,szpektor,reutt}@google.com
@@ 2 Measuring Tokenization Quality
From the very early days of NLP, models have always assumed text discretized to tokens as input Winograd (1971). For the most part, these tokens were whitespace separated words, but in the recent decade non-trivial tokenization algorithms have surfaced Mikolov et al. (2012); Sennrich et al. (2016); Ataman and Federico (2018), primarily to deal with unseen tokens without smoothing techniques or other convoluted methods Chen and Goodman (1999). The underlying reasoning behind all modern tokenization methods is that some subwords, e.g., morphemes, may carry independent information that is of value to the model even if the word as a whole is rare or unseen. Therefore, better tokenization is assumed to improve models’ performance over rare words, while also carrying computational benefits, like smaller models and the elimination of unknown tokens.
@@
Other works circle back and assess tokenization with respect to the desiderata it is supposed to serve as a stand-alone algorithm, independently from the model trained on top of it. This is done usually in addition to evaluation over downstream performance. However, most works disagree on the desiderata themselves. Many emphasize alignment to linguistically meaningful units Klein and Tsarfaty (2020); Hofmann et al. (2021, 2022); Gow-Smith et al. (2022) or to human cognitive preferences in general Beinborn and Pinter (2023). [^1] This line of works may view tokenization as a continuation of unsupervised morphemic segmentation Creutz and Lagus (2002); Virpioja et al. (2013). Others include analyses of token length and frequency Bostrom and Durrett (2020); Yehezkel and Pinter (2023), mostly in addition to the above, assuming that ideal tokenizers use longer and more frequent tokens.
The two types of tokenization evaluations, extrinsic over downstream success and intrinsic over a plethora of metrics, are usually not compared directly. They are only used to demonstrate the superiority of a specific tokenizer, and the relations between the evaluation approaches is glossed over. In this work we explicitly focus on compression as a potential intrinsic indicator of tokenization quality, as has been suggested in past works in other settings Gallé (2019); Gutierrez-Vasques et al. (2023), and check to what extent it is correlated with extrinsic downstream success. We conclude that compression is a desideratum for tokenization not only due to its theoretical virtue, expanded on in the next section, but first and foremost because it correlates with downstream performance.
@@ 3 The Role of Compression in Tokenization
But in a sense, compression-driven tokenization may be viewed as language modeling in and of itself. Consider that language models are aimed at assessing and possibly maximizing the likelihood of produced texts expressed as a product of token probabilities,
@@
where $x_{k}$ is the $k$th token in a sentence $\mathbf{x}$. Compression limits the lower bound on this product of fractions by minimizing the number of operands, i.e., minimizing the length of the sequence.
@@
$P(\mathbf{x})\approx\prod_{k}P(x_{k}|\mathbf{x}_{k-(n-1):k-1}),$
and the probability of $x_{k}$ given its context is further approximated by the number of appearances of the relevant $n$-gram in a training corpus,
$P(x_{k}|\mathbf{x}_{k-(n-1):k-1})\propto N(\mathbf{x}_{k-(n-1):k}),$
@@
We set to empirically examine our hypothesis by assessing the correlation between the tokenizer’s compression ability and the performance of language models of various sizes over a set of downstream tasks.
To explicitly control the compression ability, while fixing any other intervening factors as much as possible, we deal only with BPE tokenizers. This is in contrast with other works that compared compression across different tokenization algorithms Gallé (2019); Schmidt et al. (2024). To create BPE tokenizers with varied compression rate we recall that BPE’s maximal compression is guaranteed only over its supporting corpus. Normally, for large enough corpora, a minimal discrepancy is assumed between the character distribution in the corpus and the “true” distribution in the language. In this work however, we explicitly emphasize and expand this discrepancy by limiting the size of the support to a great extent. We will show that this intervention severely hinders the compression capacity of the tokenizer and that it also leads to deteriorating downstream performance.
@@ 4.1 English Experiments
For every tokenizer, we trained decoder-only transformer-based language models in three sizes, in terms of number of parameters: 1B, 128m and 10m. The model sizes exclude the parameters dedicated to the embedding layer, as its size may differ across tokenizers. See Appendix A for further details.
@@ 5.1 Intrinsic Evaluation
To illustrate the effect of limiting the tokenization support on the compression ability, we measured the accumulative length in tokens of the development sets of all English downstream tasks.
@@
Table 3: Spearman’s $\rho$ coefficient for rank correlation between downstream performance and the tokenizers’ support for all model sizes and tasks. Asterisks are used to denote statistically significant correlation.                                                                                                                                   * $p<0.05$, ** $p<0.01$
@@
Table 4: Pearson’s correlation coefficient between downstream performance and the tokenizers’ compression as expressed by the dev sets length (see Table 1), for all model sizes and tasks. Asterisks are used to denote statistically significant correlation.                                                                                                                                   * $p<0.05$, ** $p<0.01$
@@ 7 Conclusions
In this paper we demonstrated the importance of compression to tokenization as an intrinsic evaluation of tokenization quality that indicates the performance on extrinsic downstream tasks. We argued in favor of compression-driven tokenization from a theoretical perspective, since it may make the tokenizer act as a simple standalone language model, and we showed its correlation with downstream model success.
@@ 8 Limitations
Acknowledgements
We thank Alon Jacovi and Uri Shaham for the helpful discussions and feedback.
Appendix A Training Details
Tokenizers
were trained on the first documents in C4, for English, and mC4, for Turkish, as they are ordered in the Tensorflow Datasets repository. Therefore, the data for the less-supported tokenizers is contained in the data for the better supported ones.
[…]
In the case of char, supported by no documents at all, the tokenizer simply breaks all the words into characters and replaces any foreign character with the unknown sign.
Models
[…]